OpenStudy (wintersuntime):

Can someone be nice enough to help me with this question please ->> 5q-7= 2/3

OpenStudy (misty1212):

HI!!

OpenStudy (wintersuntime):

hi

OpenStudy (misty1212):

add 7, then divide by 5

OpenStudy (wintersuntime):

on both sides

OpenStudy (anonymous):

yes of course on both sides

OpenStudy (anonymous):

you add the 7 to get rid of the \(-7\) on the left first

OpenStudy (wintersuntime):

so if i add both of them i get 14

OpenStudy (anonymous):

then you will have \(5q\) on the left hand side to get \(q\) by itself you divide by 5

OpenStudy (anonymous):

@satellite73 Can You please help me with a question im begging

OpenStudy (anonymous):

\[5q-7=\frac{2}{3}\\ 5q=\frac{2}{3}+7\] there is no 14 in it

OpenStudy (anonymous):

5Q - 7 = 2/3 Drag 7 on other side 5q = 2/3 + 7 5q = (2+21)/7 5q = 23/7 Now drag 5 and multiply it with 7 q = 23/35 So you get Q = 0.657

OpenStudy (wintersuntime):

Isn't the answer 23/15

OpenStudy (anonymous):

Oops yeah its 23/15 my bad

OpenStudy (anonymous):

Then you will get 1.533